### v1

# Entropy-Lens: The Information Signature of Transformer Computations

Abstract
Transformer models have revolutionized fields from natural language processing to computer vision, yet their internal computational dynamics remain poorly understood—raising concerns about predictability and robustness. In this work, we introduce Entropy-Lens, a scalable, model-agnostic framework that leverages information theory to interpret frozen, off-the-shelf large-scale transformers. By quantifying the evolution of Shannon entropy within intermediate residual streams, our approach extracts computational signatures that distinguish model families, categorize task-specific prompts, and correlate with output accuracy. We further demonstrate the generality of our method by extending the analysis to vision transformers. Our results suggest that entropy-based metrics can serve as a principled tool for unveiling the inner workings of modern transformer architectures.
footnotetext: Equal contribution.
1 Introduction
Transformer-based architectures (Vaswani et al., 2023) are widely employed as state of the art models in several fields, from machine translation and search engines to DNA analysis and protein research (Devlin, 2018; Khattab and Zaharia, 2020; Ji et al., 2021; Chandra et al., 2023). Their declination in language modeling of large corpora is referred to as large language models (LLMs), and in computer vision as vision transformers (ViTs). Despite their success and ubiquity, transformers’ inner workings remain largely unknown, resulting in unpredictable behaviour (Wei et al., 2022) and reliability concerns (Schroeder and Wood-Doughty, 2025; Huang et al., 2025).
Therefore, considerate research efforts are devoted to transformer-based architectures interpretability, mostly focusing on LLMs (Nanda and Bloom, 2022; Bereska and Gavves, 2024; Elhage et al., 2021, 2022) and ViTs (Chefer et al., 2021). While exciting results have been achieved in this area, they remain limited to toy models and simplified setups, both very different from real use-case conditions. Moreover, these methods often require training a set of probes (nostalgebraist, 2020; Belrose et al., 2023) or full models on ad-hoc tasks (Nanda et al., 2023), making them architecture or even model specific and computationally expensive. These limitations restrict the scope and usability of current methodologies, rendering them unsuitable for off-the-shelf or large-scale transformer-based architectures.
Hence, we develop Entropy-Lens, a scalable framework to address these limitations. Our methodology is architecture agnostic and applicable to frozen off-the-shelf large-scale transformers. In our experiments, we consider both LLMs, including Llama (Touvron et al., 2023), Gemma (Team et al., 2024), GPT (Radford, 2018) up to 9B parameters, and ViTs (Wu et al., 2020) and data-efficient image transformers (DeiTs) (Touvron et al., 2021; Wightman, 2019). In particular, we analyze the evolution of the generated tokens’ Shannon entropy after each intermediate block in the residual stream, as described in Section 4. In Section 5, we show that these quantities constitute an information-theoretic signature of which and how the computation is performed. Finally, our work paves the way for several potential research directions involving information theory for LLM interpretability, as outlined in Section 6.
Our contributions are as follows:
  1. 1.
    We develop a scalable model agnostic methodology grounded in information theory for frozen off-the-shelf large-scale transformer architectures interpretability (Section 4).
  2. 2.
    We analyze the Shannon entropy of LLMs’ generated tokens’ intermediate predictions, showing that these quantities (1) identify the model family that generated them (2) identify the ‘task type’ of the prompt (3) correlate with the correctness of LLM generated answers to multiple choice questions (Section 5).
  3. 3.
    We demonstrate that the same framework is adaptable to domains outside of language modeling. In particular, we apply it to computer vision to show that the Shannon entropy of ViTs’ intermediate predictions identify the model family that generated them (Section 5), similarly to the LLMs case.
2 Related Work
Lenses in LLMs
Mechanistic interpretability (Bereska and Gavves, 2024) aims to provide a precise description and prediction of transformer-based computations. Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream. For example, logit-lens (nostalgebraist, 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space. tuned-lens (Belrose et al., 2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function. Building on the Transformer-Lens library (Nanda and Bloom, 2022), we propose Entropy-Lens, which employ logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
Transformers’ Circuits
Another approach to mechanistic interpretability aims to identify and understand the specific sub-computations, or circuits, in a neural network (Saphra and Wiegreffe, 2024). Olah et al. (2020) pioneered this approach, introducing the concept of circuits and demonstrating their existence in small models through manual analysis. In transformers, circuits are hypothesized to act as agents that read from and write to the residual stream, which acts as a form of memory. This has been demonstrated in a simplified transformer model composed only of attention blocks (without MLPs) (Elhage et al., 2021). However, there is evidence that full transformers exhibit similar behavior. This evidence—linked to the concept of superposition, the idea that models can represent more features than the available dimensions by compressing multiple features into one (Elhage et al., 2022)—is supported by studies on sparse autoencoders, which demonstrate the ability to decompose representations into simpler components (Bricken et al., 2023). Conmy et al. (2023a) developed a toolkit to facilitate mechanistic interpretability, offering techniques like activation patching and weight factorization. Building upon this, Conmy et al. (2023b) explored automated circuit discovery methods, addressing the challenge of scaling analysis to larger models. These works collectively emphasize the importance of understanding the concrete computational steps within LLMs, moving beyond superficial observations to reveal the underlying mechanisms.
Information Theory in Transformers
Information theory has been studied both in connection to the training phase of LLMs and their interpretability. For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al., 2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al., 2024). Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton et al., 2024). Our work, instead, shifts the focus on the vocabulary’s natural domain. Through Entropy-Lens, we use information theory to study the evolution of entropy of the intermediate layers’ decoded logits.
3 Background
3.1 Information Theory
The main information-theoretic quantity used in our study is entropy. Given a discrete [^1] Although entropy can be naturally extended to the continuous case with probability density functions, we restrict ourselves to the discrete case as it is the most relevant to our study. random variable $X$ with outcomes $x_{i}$ and probability mass function $p$, the Shannon entropy $H$ of $X$ is defined as
$H(X)=-\sum_{i}p(x_{i})\log p(x_{i})=\mathbb{E}[-\log p(X)]$ (1)
Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’. These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible event $A$ and $B$. The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined.
Next, we study the entropy of vocabulary predictions—a quantity that is maximal when the prediction assigns equal probability to all tokens, minimal when it assigns zero probability to all but one token, and takes intermediate values when probability is distributed across multiple tokens, consistent with the previously mentioned properties.
3.2 The Transformer
Architecture
Figure 1: Structure of a generic Transformer block.
The transformer (Vaswani et al., 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021). Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 2. While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich, 2019), attention, and multi layer perceptrons (MLPs), as shown in Fig. 1. Residual connections, instead, sum the output of the layer $i-1$ to the output of the layer $i$.
Inside a single transformer block, the information flows both horizontally and vertically. The former, enabled by the attention mechanism, allows the token representations to interact with each other. In a language modelling task, for example, this is useful to identify which parts of the input sequence—the sentence prompt—should influence the next token prediction and quantify by how much. The latter vertical information flow allows the representation to evolve and encode different meanings or concepts. Usually, the dimension of the latent space is the same for each block in the transformer. The embedding spaces where these computations take place are generally called the residual stream.
Computation schema
LLMs are trained to simply predict the next token in a sentence. That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$. Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, sometimes set to $E^{\top}$, tying the embedding weights and the decoder. Finally, the logits—the output of $D$—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary. We summarize this operation with the function $W:=\operatorname{softmax}\circ\,D$.
In formal terms, information processing can be expressed using the encoder, decoder, Transformer block $f$, and residual connection:
$x_{j}^{0}$ $=E(t_{j})$ (2)
$x_{j}^{i}$ $=f^{i}(x_{j}^{i-1})+x_{j}^{i-1}$ (3)
$y_{j}^{i}$ $=W(x_{j}^{i})$ (4)
where $j\in\{1,\dots,N\}$ ranges over the number of tokens in the prompt and $i\in\{0,\dots,L\}$ ranges over the number of layers. Hence, $x_{j}^{i}$ represents the activations of token $t_{j}$ after layer $i$.
3.2.1 Instruct Models
Training a Large Language Model (LLM) requires vast amounts of data and is generally divided into multiple phases.
Pretraining: The model is exposed to large datasets through self-supervised tasks, such as next-token prediction or similar variants. This phase helps the model learn a broad range of general knowledge.
Fine-tuning: This phase teaches the model to generate more useful and coherent responses. Two main strategies are used: Chat: The model is trained on structured conversations between a user and the model, with clearly defined roles. Instruct: The model learns from simple commands, without a predefined dialogue structure.
RLHF (optional): Some models undergo Reinforcement Learning from Human Feedback (RLHF) to further refine their responses based on human preferences.
For our experiments, we used off-the-shelf models without RLHF to analyze information processing in a less biased transformer version. We also focused on Instruct models instead of Chat models for two reasons: 1. the Instruct strategy aligns better with our experimental setup 2. Instruct models are more flexible and often preferred for practical applications.
4 Method
The aim of our framework is to find and characterize the information-theoretic signature of transformer computations. Entropy-Lens’s pipeline comprises three steps and is described in Figure 2.
Notation
We denote the input sentence comprising tokens $t_{1},\dots,t_{N}$ by $S=(t_{i})_{i=1}^{N}$. Then, $x_{j}^{i}$ denotes the activations of the token $t_{j}$ after block $i$ for $j\in\{1,\dots,N\}$ and $i\in\{1,\dots,L\}$. Since our analysis focuses on the logits extracted from the intermediate layers of the transformer, it will be useful to distinguish between normalized and unnormalized logits. We define $W:=\operatorname{softmax}\circ\,D$ and $y_{j}^{i}:=W(x_{j}^{i})$ the normalized logits of the token $t_{j}$’s activations after layer $i$.
The core of our methodology is to analyze the entropy of the generated tokens’ intermediate representations $y_{j}^{i}$. These vectors are probability distributions, as they are the output of a $\operatorname{softmax}$. To obtain a single quantity that summarizes the information they contain, we compute their entropy $H(y_{j}^{i})$. For one generated token, we can consider the entropy of all of its intermediate predictions $H(y_{j}^{i})$ for $i\in\{1,\dots,L\}$. This leads us to the definition of entropy profile:
Defintion 1 (Entropy profile)
Let $h_{j}^{i}=H(y_{j}^{i})$ be the entropy of the intermediate representation of token $t_{j}$ after block $i$ and residual connection. The entropy profile of the next generated token is defined as
$h_{N}=\bigoplus_{i}h_{N}^{i}$ (5)
where $\bigoplus$ denotes any aggregation function.
In our experiments, we set $\bigoplus$ to be concatenation, so that $h_{N}=(h_{N}^{1},\dots,h_{N}^{L})^{\top}$, but other choices are possible. The extraction of entropy profiles is the step 1 of our pipeline.
Then, we fix the number of tokens that the LLM is required to generate, $T$ and repeat the same procedure for each of them, leading us to the next definition:
Defintion 2 (Aggregated entropy profile)
Let $h_{N+t}$ be the entropy profiles according to Definition 1 for $t\in\{0,\dots,T-1\}$, i.e.  the entropy profile of each token generated sequentially by a transformer. The aggregated entropy profile of the next $T$ generated tokens is defined as
$h_{[N:T]}=\bigotimes_{t=0}^{T-1}h_{N+t}$ (6)
where $\bigotimes$ denotes any aggregation function.
Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ defined in Definition 1. In our experiments, we set both of them to be concatenation, so that $h_{[N:T]}$ is the matrix with $h_{N+t}$ as columns, that is $(h_{[N:T]})_{t}^{i}=h_{N+t}^{i}$ for $i\in\{1,\dots,L\}$ and $t\in\{0,\dots,T-1\}$. The aggregation of entropy profiles is the step 2 of our framework.
The last step of our framework is classification, where we feed the aggregated entropy profile to a classifier $\mathcal{C}$ to determine whether it contains sufficient information to identify a particular ‘entity’.
In our experiments, we examine whether aggregated entropy profiles identify model family (Section 5.1.1), task type (Section 5.1.2), and correct and wrong answers to multiple choice questions (Section 5.1.3) in LLMs. Additionally, we apply the same pipeline to ViTs showing the flexibility of the proposed methodology (Section 5.2). In our experiments, we take $\mathcal{C}$ to be a k-NN classifier. Classification of the aggregated entropy profiles is the step 3 our framework.
Figure 2: Entropy-Lens’s pipeline. (Left) A diagram representing a transformer architecture: hidden representations are converted into intermediate predictions with $W$ before calculating their entropy $H$. (Right) A diagram representing our framework: step 1: entropy profile extraction, step 2: entropy profile aggregation and step 3: classification.
5 Experiments
The goal of our experiments is to demonstrate the effectiveness of entropy profiles in distinguishing computational signatures across various models and scenarios. In the first part, we explore how entropy profiles can differentiate between LLM families and characterize distinct tasks. In the second part, we apply our framework to the Vision Transformer architecture to assess whether insights from the LLM experiments can be similarly recovered.
5.1 Entropy-Lens for Large Language Models
The experiments on LLMs focus on three key aspects. First, we demonstrate that entropy profiles are indicative of model families, with distinctions becoming more pronounced as model size increases. Second, we investigate whether the entropy profile of a model alone can be used to classify the task it is performing. Finally, we extend this analysis to assess whether entropy profiles can distinguish between correct and incorrect task execution.
5.1.1 Entropy profiles identify model families
We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 9 models from 3 different families (GPT, Gemma and LLama) with parameter counts ranging from 100M to 9B.
We compute the mean of the entropy profiles obtained from 64 generated tokens, obtained with the prompt ‘The concept of entropy, a brief essay:’, as shown in Figure 3. We observe that the profiles relate uniquely to the model family, rather than a particular model, independently of its size.
The GPT model class starts with high vocabulary entropy in the early layers, indicating a wide range of possible response tokens. Then, entropy gradually decreases—more smoothly than in other classes—leading to a low-entropy state, where the model narrows down to a small set of possible response tokens.
The Gemma model class, on the other hand, starts with low entropy in the very first layers, then rises to higher entropy in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
The Llama model class follows a similar pattern, but with a steeper rise, resulting in a higher entropy value maintained over a larger range of intermediate layers.
We observe that the equivalence between models of the same family but different sizes holds when looking at the entropy trend not as a function of the absolute layer index, but rather as the relative layer position within the model.
We conjecture that high entropy phases, whether in the early or intermediate layers, allow the model to explore more possibilities in its response, similarly to how temperature helps avoid getting stuck in local minima in optimization. Then, at the moment of selection, the distribution is ‘cooled down’, forcing the output to be limited to a few possible tokens.
Figure 3: Average entropy profiles over 64 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
5.1.2 Entropy profiles identify task types
We verify whether the entropy profiles can identify task types examining generative (continue a text), syntactic (count the number of words in a text), and semantic (extract the subject or moral of a text) tasks.
We do this with the TinyStories dataset (Eldan and Li, 2023). For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix A Table 3, and a story from TinyStories. These templates are:
  • Base, of the form task prompt + story
  • Reversed, of the form story + task prompt
  • Scrambled, of the form task prompt + scrambled story or scrambled story + task prompt, at random. A scrambled story is a ‘story’ obtained by randomly shuffling the words in a given story from TinyStories.
Note that, for a robust evaluation, we also use 2 possible task prompt variations, as per Table 3.
We generate 800 prompts per task type, $1/3$ of them with the base template, $1/3$ with the reversed template, and $1/3$ with the scrambled templates, for a total of 2400 prompts. We then apply our pipeline from Section 4 to classify the aggregated entropy profiles of these prompts against their task type using a k-NN classifier. The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest). Table 1 shows the results obtained for 6 models with parameter counts ranging from 1B to 9B. Figure 4 shows the average entropy profiles per task type.
We observe that all k-NN classifiers (i.e. one for each LLM) achieve high accuracy in distinguishing entropy profiles, with a trend toward improved performance for larger models.
Figure 4: Average entropy profiles with shaded standard deviation for different task types: generative (continuation_prompt 1 and 2), syntactic (counting_prompt 1 and 2), and semantic (semantic_prompt 1 and 2). These tasks are induced with the prompts described in Appendix A. Left: Llama-3.2-Insruct. Right: Gemma-2-Instruct
Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. The standard deviation is calculated on the 10-fold cross-validation splits.
Model Model Size k-NN ROC-AUC
Gemma-2-it 2.1B 97.66 ± 0.47
Gemma-2-it 8.9B 98.38 ± 0.50
Llama-3.2-Instruct 1 B 94.94 ± 0.79
Llama-3.2-Instruct 3 B 94.77 ± 0.93
Llama-3-Instruct 8 B 96.10 ± 0.67
Phi-3 3.6B 97.07 ± 0.87
5.1.3 Entropy profiles identify correct task execution
We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset Hendrycks et al. (2021).
MMLU consists of multiple-choice questions across 57 subjects, ranging from history and physics to law, mathematics, and medicine. The difficulty levels span from elementary to professional, making it a benchmark for evaluating language models on specialized domains. Each dataset entry contains: a question string, four answer choices and a label indicating the correct answer.
We evaluated two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 4 in Appendix A):
  • Base: A minimal version containing the topic, the question, and multiple-choice answers.
  • Instruct: A version with a brief explanation that it’s a multiple-choice test where only one option should be selected.
  • Humble: A version that also instructs the model to pick a completely random option if it doesn’t know the answer.
Then, we applied our pipeline to extract and aggregate and the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer. We train a k-NN classifier for each LLM and validate it using 10-fold cross-validation. We also conducted a t-test to compare our classifier to a dummy model. This dummy model generates predictions randomly, sampled from a distribution that reflects the proportion of correct and incorrect answers produced by the LLM, ensuring robustness against class imbalance. The results reject the null hypothesis with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (Table 2).
We observe that the instruct and humble prompts improve Llama’s average accuracy, while for Gemma, this is only true for the instruct prompt. Additionally, in Llama, the model’s higher accuracy seems to be partially linked to greater difficulty in distinguishing correct from incorrect entropy profiles, though more rigorous analysis is needed to confirm this. In Gemma, however, this claim is harder to support.
Figure 5: Entropy profiles for ViT model families.
5.2 Entropy-Lens for Vision Transformers
Finally, to demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
Using 20 classes from ImageNet-1K (Russakovsky et al., 2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 5. We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers. This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
Focusing on computer vision models, we note that while ViT and DeiT families exhibit qualitatively similar trends, they differ quantitatively—ViTs start with higher entropy values, making them easily distinguishable from DeiTs.
Notably, the only profile that stands out is that of ViT Large (with $\sim 300$M parameters), compared to the other models analyzed in this section, which have $\leq 86$M parameters.
For ViT Large, entropy decreases more smoothly, appearing not only as a better approximation of the sharp drop seen in smaller models but possibly following a different behavior entirely, with the entropy decline starting earlier.
We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between $87$M and $307$M parameters.
Table 2: Performance comparison of different models and prompt types on MMLU accuracy and k-NN AUC-ROC score.
Model Prompt Type LLM Accuracy k-NN AUC-ROC
Llama-3.2-3B-Instruct Base 50.89 73.61 ± 1.52
Humble 58.51 69.90 ± 1.06
Instruct 60.62 67.23 ± 1.62
Gemma-2-2B-it Base 56.10 71.88 ± 1.63
Humble 54.71 72.78 ± 1.15
Instruct 56.38 68.36 ± 1.23
6 Conclusions
In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory. In Section 5.1.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family. In Section 5.2 we conduct similar experiments on vision transformers, demonstrating the wide applicability of our framework. Additionally, in Section 5.1.2 we showed that the same entropy profiles can be used to identify the ‘task type’ in LLMs, and in Section 5.1.3 we used them to distinguish between correct and wrong LLM generated answers. Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
Appendix A Evaluation Details
Table 3: Prompt templates used for TinyStories tasks.
Task Type Task prompt
Generative How can the story be continued?
Which could be a continuation of the story?
Syntactic How many words are in the story?
Count the number of words in the story.
Semantic What is the main idea of the story?
Who is the subject of the story?
Table 4: Prompt templates used for the MMLU dataset.
Prompt Type Prompt
Base
Subject: {subject}
Question: {question}
Choices:
A. {option_1}
B. {option_2}
C. {option_3}
D. {option_4}
Answer:
Instruct
The following is a multiple-choice question about {subject}. Reply only with the correct option.
Question: {question}
Choices:
A. {option_1}
B. {option_2}
C. {option_3}
D. {option_4}
Answer:
Humble
The following is a multiple-choice question about {subject}. Reply only with the correct option.
If you are unsure about the answer, reply with a completely random option.
Question: {question}
Choices:
A. {option_1}
B. {option_2}
C. {option_3}
D. {option_4}
Answer: